First estimate the answer. Then divide to find the exact answer. simply all answers.
1 and 1/5 divided by 3 and 3/5

Answer by checkley79(3341)

1 and 1/5 divided by 3 and 3/5
1/3=.3333333 EST.
1.2/3.6=1/3 ACTUAL.